OpenStudy (anonymous):

simplify by using a double angle or a half angle formula 1-cos 6x ___________ sin 6x

OpenStudy (ddcamp):

cos(6x) = 1 - 2sin²(3x), so the numerator can be simplified to 2sin²(3x) sin(6x) = 2sin(3u)cos(3u) So we have: \[\frac{ 2\sin^2(3x) }{ 2\sin(3x)\cos(3x) } = \frac{ \sin(3x) }{ \cos(3x) } = \tan(3x)\]

OpenStudy (anonymous):

yeah tan3x ans

OpenStudy (anonymous):

thx both

OpenStudy (anonymous):

tan 3x